# Address-addressable read AI artificial intelligence

### v1

@@ 4.1 Semantic Reasoning
To achieve this, we need to represent the knowledge encoded in the premises and write a small programme that employs calls to our Views-based GDB to work through the necessary logical steps for deriving the conclusion. For the representation let us choose a straightforward scheme: First, a chain to represent the object we are attempting to make an inference about (“This”) including a linknode with primIDs “species” and “cat” (just as we have done in Figure 3b). Then, a chain for the general concept of a “cat” featuring a linknode whose primIDs are the pair (“family”, “Felidae”) plus any other relevant chains. These are illustrated in Figure 9a. Next, comes the programme. This can be code running on any host machine with access to a Views-based GDB. We will focus on the calls to the GDB here before presenting the full pseudocode in Algorithm 1. Throughout this example, we write the database call code in a form that closely resembles Python for simplicity and familiarity. We also omit trivial checks such as verifying that our CAR/AAR operations return valid results to focus on the logic of the algorithm rather than the minutiae.
@@ 6 Conclusions
In this paper, we have proposed a GDB model named Views, which supports the notion of graphs with infinitely recursive labellability (whereby edges and vertices, and their properties, and their properties ad infinitum can be labelled) in an intuitive manner, and converts the infinitely recursively labellable graphs into linked list structures for hardware-friendly data storage and graph traversal. Furthermore, we have laid the foundations of a natural hardware implementation of our model and provided a fundamental instruction set architecture (ISA) that can operate on the Views data structure, including massively-parallel content-addressable read operations. This hints towards the potential for tremendous acceleration of graph analytics workloads via parallelisation. Next, we showed examples of how the structure can be used to carry out example reasoning tasks and how it may be naturally extended to accommodate the operational requirements of a non-trivial example cognitive model, Copycat, which uses bespoke features such as its “slippage” mechanism. This shows promise towards eventually underpinning key cognitive tasks such as semantic reasoning, logical deduction, and cognitive processes such as analogy. Finally, we have illustrated how the approach is versatile and highly compatible with (i.e. “admits mappings from”) a range of conventional knowledge graph representations, which indicates a relatively low barrier to the translation from well-established representations into Views. We hope that our proposed data structure and the corresponding hardware implementation outline act as an extra bridge between the computer science & AI community and the hardware design community, inspiring further innovation in hardware-aware AI long into the future.

### v2

@@ 4.1 Semantic Reasoning
To achieve this, we need to represent the knowledge encoded in the premises and write a small programme that employs calls to our Views-based GDB to work through the necessary logical steps for deriving the conclusion. For the representation let us choose a straightforward scheme: First, a chain to represent the object we are attempting to make an inference about (“This”) including a linknode with primIDs “species” and “cat” (just as we have done in Figure 3b). Then, a chain for the general concept of a “cat” featuring a linknode whose primIDs are the pair (“family”, “Felidae”) plus any other relevant chains. These are illustrated in Figure 10a.
Next, comes the programme. This can be code running on any host machine with access to a Views-based GDB. We will focus on the calls to the GDB here before presenting the full pseudocode in Algorithm 1. Throughout this example, we write the database call code in a form that closely resembles Python for simplicity and familiarity. We also omit trivial checks such as verifying that our CAR/AAR operations return valid results to focus on the logic of the algorithm rather than the minutiae.
@@ 6 Conclusions
In this paper, we have proposed a GDB model named Views, which supports the notion of graphs with infinitely recursive labellability (whereby edges and vertices, and their properties, and their properties ad infinitum can be labelled) in an intuitive manner, and converts the infinitely recursively labellable graphs into linked list structures for hardware-friendly data storage and graph traversal. Furthermore, we have laid the foundations of a natural hardware implementation of our model and provided a fundamental instruction set architecture (ISA) that can operate on the Views data structure, including massively-parallel content-addressable read operations. Next, we evaluated this hardware implementation’s storage efficiency across existing RDF and LPG-based GDB implementations revealing an advantage in storage efficiency and scalability stemming from the combination of a uniform data structure, a linked-list organisation and tight co-design with the corresponding hardware. This hints towards the potential for tremendous acceleration of graph analytics workloads via parallelisation. Next, we showed examples of how the structure can be used to carry out example reasoning tasks and how it may be naturally extended to accommodate the operational requirements of a non-trivial example cognitive model, Copycat, which uses bespoke features such as its “slippage” mechanism. This shows promise towards eventually underpinning key cognitive tasks such as semantic reasoning, logical deduction, and cognitive processes such as analogy. Finally, we have illustrated how the approach is versatile and highly compatible with (i.e. “admits mappings from”) a range of conventional knowledge graph representations, which indicates a relatively low barrier to the translation from well-established representations into Views. We hope that our proposed data structure and the corresponding hardware implementation outline act as an extra bridge between the computer science & AI community and the hardware design community, inspiring further innovation in hardware-aware AI long into the future.
Acknowledgements
This study was supported by the Engineering and Physical Sciences Research Council (EPSRC) under Grant EP/V008242/2 and Scottish Enterprise High Growth Spinout Programme (HGSP) Atlas, funding reference number PS7305129O.